# Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade and with a quantized radiation field studied by the Jaynes-Cummings Model

### v1

@@ Abstract
The interaction between atoms and a quantized radiation field is fundamentally important in quantum optics and quantum information science. Due to their unusual properties, Rydberg atoms are promising building blocks for two-qubit gates and atom-light quantum interfaces, exploiting the Rydberg blockade interaction which prevents two atoms at close distance (< 10 $\mu$m) from being simultaneously excited to Rydberg states. Recently, this effect was used to engineer quantum processors based on arrays of interacting Rydberg atoms illuminated by Raman lasers. Motivated by these experiments, we extend the Jaynes-Cummings model to study the interaction between two Rydberg atoms interacting by the Rydberg blockade and a quantized radiation field. We find a rich variety of Rabi oscillations and entanglement between the atoms and the radiation field as a function of initial conditions and interaction time, which may be used to obtain atom-light quantum interfaces as components for future long-distance quantum communication.
@@ 1 Introduction
Among the striking properties of quantized light-matter interactions described by the JCM are entanglement between the atoms(s) and the quantized radiation [20, 27, 5, 15, 21, 28, 29, 30], and the collapse and revival of periodic populations of atomic states (Rabi oscillations) for quantum coherent states of the field [31, 13]. These phenomena are entirely due to the quantum nature of the field and do not occur in a semiclassical treatment; indeed, the first experimental verification of field quantization according to the JCM [3, 4] was the observation of the collapse-revival phenomenon predicted in [31].
@@
|$\ket{R}=$ (2)
$\frac{1}{\sqrt{N}}\sum\limits_{j=1}^{N}e^{i{\bf k}\cdot{\bf x}_{j}}\ket{g_{1},\ldots,g_{j-1},r_{j},g_{j+1},\ldots,g_{N}}$
in which one Rydberg excitation is coherently shared between all atoms in the ensemble [57, 58, 59]. In (2) |$\ket{g_{j}}$ and |$\ket{r_{j}}$ denote the ground state and a Rydberg state of atom $j$, respectively, ${\bf x}_{j}$ is the position of atom $j$, and $\bf{k}$ is the wave vector of the light mode. Under conditions of a perfect Rydberg blockade any state with two or more atoms excited to |$\ket{r}$ is inaccessible to the system; in this case, the ensemble of $N$ atoms is considered to be an effective two-level Rydberg “superatom” with ground state |$\ket{G}$ and excited state |$\ket{R}$. However, of course, states other than |$\ket{G}$ and |$\ket{R}$ are possible, depending on the initial conditions of the system; for example, the states |$\ket{r,g}$ and |$\ket{g,r}$ can occur in the case of two atoms (Section 3.3). The collective nature of the excitation in |$\ket{R}$ in (2) enhances the coupling of light to a Rydberg superatom by a factor of $\sqrt{N}$; as a result, Rydberg systems with $N\gg 1$ (several hundreds to thousands) atoms have attracted tremendous interest in atomic physics and quantum optics [57, 58, 41, 59]. The Rydberg blockade effect combined with the strong coupling between Rydberg superatoms and light have led to numerous applications of Rydberg atoms in quantum optics and quantum information science, reviewed in [40, 60, 42].
@@
Motivated by these experiments, in this paper we extend the Jaynes-Cummings model to study two interacting Rydberg atoms under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible (Figure 1). Our motivation is to better understand the interplay between interacting Rydberg atoms and a quantized radiation field in view of the physical implementation of quantum state-transfer between atoms and photons for future quantum information processing [33, 35, 36, 28, 29, 30] and photon-photon quantum gates [66, 67, 68, 41]. In Section 2 we formulate the Hamiltonian of the model and discuss its eigenstates and eigenvalues. We then discuss the dynamics of the system for different initial conditions, including Rabi oscillations and entanglement, for number (Fock) states of the radiation field (Section 3) and a quantum coherent field (Section 4). We conclude in Section 5. Detailed derivations of some results discussed in the main text are left to Appendices A - C.
@@ 2 Hamiltonian and eigenstates
In this section, we extend the Jaynes-Cummings model to two Rydberg atoms interacting by the Rydberg blockade. We consider two two-level atoms $a$ and $b$ with ground states |$\ket{g_{a}}$, |$\ket{g_{b}}$ and Rydberg states |$\ket{r_{a}}$, |$\ket{r_{b}}$  [^2] When there is no danger of confusion we omit the labels $a$ and $b$ and denote the ground and Rydberg states for both atoms by |$\ket{g}$ and |$\ket{r}$, respectively., respectively, interacting with a single mode of a quantized radiation field ($f$) with (angular) frequency $\omega_{f}$  [^3] If not stated otherwise, we use the term “frequency” for the angular frequency in units of $\text{rad}\cdot\text{s}^{-1}$.. We assume that the unperturbed transition frequencies $\omega_{0}=\omega_{r}-\omega_{g}$ of the atomic transition |$\ket{g}\leftrightarrow\ket{r}$ of both atoms are equal. The combined system of the two atoms and the quantized field is spanned by product states of the form
@@
=|$=\ket{\psi_{a},\psi_{b},\phi_{f}}\,.$
We assume that the Rydberg blockade between the atoms is in full effect, so that the doubly excited state, |$\ket{r,r,\phi_{f}}$, is not accessible to the system. Thus, accessible states are spanned by states of the form |$\ket{g,g,n}$, |$\ket{g,r,n}$, and |$\ket{r,g,n}$, with the photon number (Fock) states |$\ket{n}$, $n=0,1,2,\ldots$, of the field. We describe the interaction between the atoms and the quantized field by a Jaynes-Cummings-type Hamiltonian in the rotating wave approximation,  [^4] To simplify notation, in (4) we use the $\otimes$ symbol only to separate operators acting on atoms $a$ and $b$, respectively, but not for the field operators $\hat{a}$, $\hat{a}^{{\dagger}}$.
$\hat{H}$ $=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{0}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (4)
$\,+\hbar\lambda_{a}\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\right)\otimes{\hat{P}}_{g}^{b}$
$\,+{\hat{P}}_{g}^{a}\otimes\hbar\lambda_{b}\left({\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,.$
The first term in (4) is the energy of the field with the photon number operator ${\hat{n}}={\hat{a}}^{\dagger}{\hat{a}}$, where ${\hat{a}}^{\dagger}$, $\hat{a}$ are bosonic creation and annihilation operators and $\hbar=h/\left(2\pi\right)$ is the reduced Planck constant. The second term in (4) is the energy of the unperturbed atoms $a$, $b$ (upper labels) expressed in terms of the Pauli operators ${\hat{\sigma}}_{z}=\mathinner{|{g}\rangle\langle{g}|}-\mathinner{|{r}\rangle\langle{r}|}$, where the zero of energy is chosen as the average of the energies of |$\ket{g}$ and |$\ket{r}$. The interaction terms in the second and third lines of (4) describe processes in which either atom $a$ or atom $b$ is excited from |$\ket{g}$ to |$\ket{r}$ by absorption of one photon, or decays from |$\ket{r}$ to |$\ket{g}$ by emission of one photon, with coupling constants $\lambda_{a}$, $\lambda_{b}$ for atoms $a$, $b$, respectively; ${\hat{\sigma}}_{+}=\mathinner{|{r}\rangle\langle{g}|}$ and ${\hat{\sigma}}_{-}=\mathinner{|{g}\rangle\langle{r}|}$ are raising and lowering Pauli operators for atoms $a$, $b$ (upper labels in (4)). In the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ the Pauli operators are represented by $2\times 2$ matrices  [^5] Note that in the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ (in this order) the raising and lowering Pauli matrices $\sigma_{+}$, $\sigma_{-}$ in (5) are defined the opposite way compared to raising and lowering Pauli matrices for spin-$1/2$ particles.
@@
Since the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system due to the Rydberg blockade, excitation and decay of one atom can only occur when the other atom is in its ground state |$\ket{g}$; this effect is incorporated in the Hamiltonian (4) by the projection operators ${\hat{P}}_{g}^{a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of atoms $a$ and $b$, respectively.
@@
$\hat{H}$ $=\hbar\omega_{f}\hat{N}+\frac{1}{2}\hbar\Delta\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)$ (9)
$\,+\hbar\lambda\Big{[}\Big{(}{\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\Big{)}\otimes{\hat{P}}_{g}^{b}$
$\,+{\hat{P}}_{g}^{a}\otimes\Big{(}{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\Big{)}\Big{]}\,,$
@@
|$\ket{\psi_{1}^{(n)}}:=\ket{r,g,n}\,,\,\,\ket{\psi_{2}^{(n)}}:=\ket{g,r,n}\,,\,\,$ (11)
|$\ket{\psi_{3}^{(n)}}:=\ket{g,g,n+1}\,\,.$
This implies that the total Hilbert space $\mathbb{H}$ of the system decouples into subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$, and $\hat{H}$ in (9) can be expressed as a direct sum ${\hat{H}}=\mathop{\oplus}\limits_{n=0}^{\infty}{\hat{H}}^{(n)}$ where the action of each component ${\hat{H}}^{(n)}$ is confined to the subspace $\mathbb{H}_{n}$. In the basis {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ the Hamiltonian ${\hat{H}}^{(n)}$ is represented by a $3\times 3$ matrix
@@
|$\ket{\text{asym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)$ (17)
$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)\,\,.$
@@
is the $\Delta$-dependent Rabi frequency for the two-atom system, which is different from the Rabi frequency $\Omega_{n}(\Delta)$ for one atom [1, 13] by an extra factor of 2 in front of the term $\lambda^{2}(n+1)$ inside the square root. The normalization constants in (16) are $N_{\pm}=\sqrt{\left(\Omega_{n}\pm\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$.
@@
Geometrically, $R$ in (22) describes a rotation of the eigenvectors |±$\ket{\pm,n}$ by the angle $\varphi_{n}(\Delta)$ about the fixed axis given by the $\Delta$-independent eigenvector |$\ket{\text{asym},n}$ (Figure 3).
@@
|$\ket{\psi_{i}^{(n)}}=R_{\text{asym},\,i}^{T}\ket{\text{asym},n}$ (25)
$\quad+R_{+,\,i}^{T}\ket{+,n}+R_{-,\,i}^{T}\ket{-,n}\,,\,\,i=1,2,3\,\,,$
@@
|±$\ket{\pm,n}=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{\psi_{3}^{(n)}}\right)$ (33)
$\quad=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{g,g,n+1}\right)\,,$
@@
|$\ket{\text{sym},n}$ $:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)$ (34)
$=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)\,\,.$

### v2

# Rabi oscillations and entanglement between two atoms interacting by the Rydberg blockade studied by the Jaynes-Cummings Model

@@ Abstract
The interaction between atoms and a quantized radiation field is fundamentally important in quantum optics and quantum information science. Due to their unusual properties, Rydberg atoms are promising building blocks for two-qubit gates and atom-light quantum interfaces, exploiting the Rydberg blockade interaction which prevents two atoms at close distance ($<$ 10 $\mu$m) from being simultaneously excited to Rydberg states. Recently, this effect was used to engineer quantum processors based on arrays of interacting Rydberg atoms illuminated by Raman lasers. Motivated by these experiments, we extend the Jaynes-Cummings model to study the interaction between two Rydberg atoms interacting by the Rydberg blockade and a quantized radiation field. We consider both number (Fock) states of the field and single-mode quantum coherent states. In particular, we discuss different types of entanglements between various components of the total system consisting of the two Rydberg-interacting atoms and coherent states of the field, and show that the behavior is significantly different compared to a system with non-interacting atoms corresponding to the two-atom Tavis-Cummings model. Our results are relevant in view of atom-light quantum interfaces as components for future long-distance quantum communication.
keywords:
Jaynes-Cummings model , cavity QED , Rydberg atoms , light-matter interactions
journal: Physics Open
[…]
[add]organization=Department of Physics and Astronomy, University of Texas Rio Grande Valley, addressline=1 W University Blvd, city=Brownsville, postcode=TX 78520, country=USA
@@ 1 Introduction
Among the striking properties of quantized light-matter interactions described by the JCM are entanglement between the atoms(s) and the quantized radiation [20, 27, 28, 5, 15, 21, 29, 30, 31], and the collapse and revival of periodic populations of atomic states (Rabi oscillations) for quantum coherent states of the field [32, 28, 13]. These phenomena are entirely due to the quantum nature of the field and do not occur in a semiclassical treatment; indeed, the first experimental verification of field quantization according to the JCM [3, 4] was the observation of the collapse-revival phenomenon predicted in [32]. We note that in the Dicke model and Tavis-Cummings model, the atoms do not interact directly with one another but may become entangled via their coupling to a common cavity field [20, 27, 28].
@@
|$\ket{R}=\frac{1}{\sqrt{N}}\sum\limits_{j=1}^{N}e^{i{\bf k}\cdot{\bf x}_{j}}\ket{g_{1},\ldots,g_{j-1},r_{j},g_{j+1},\ldots,g_{N}}$ (2)
in which one Rydberg excitation is coherently shared between all atoms in the ensemble [58, 59, 60, 61]. In (2) |$\ket{g_{j}}$ and |$\ket{r_{j}}$ denote the ground state and a Rydberg state of atom $j$, respectively, ${\bf x}_{j}$ is the position of atom $j$, and $\bf{k}$ is the wave vector of the light mode. Under conditions of a perfect Rydberg blockade any state with two or more atoms excited to |$\ket{r}$ is inaccessible to the system; in this case, the ensemble of $N$ atoms is considered to be an effective two-level Rydberg “superatom” with ground state |$\ket{G}$ and excited state |$\ket{R}$. However, of course, states other than |$\ket{G}$ and |$\ket{R}$ are possible, depending on the initial conditions of the system; for example, the states |$\ket{r,g}$ and |$\ket{g,r}$ can occur in the case of two atoms (Section 3.3). The collective nature of the excitation in |$\ket{R}$ in (2) enhances the coupling of light to a Rydberg superatom by a factor of $\sqrt{N}$; as a result, Rydberg systems with $N\gg 1$ (several hundreds to thousands) atoms have attracted tremendous interest in atomic physics and quantum optics [58, 59, 60, 42, 61]. The Rydberg blockade effect combined with the strong coupling between Rydberg superatoms and light have led to numerous applications of Rydberg atoms in quantum optics and quantum information science, reviewed in [41, 62, 43].
@@
Motivated by these experiments, in this paper we extend the Jaynes-Cummings model to study two interacting Rydberg atoms under perfect Rydberg blockade conditions, so that the doubly excited state |$\ket{r,r}$ is not accessible (Figure 1). This direct interaction between the atoms makes the model fundamentally different from the Dicke model and Tavis-Cummings model, in which the atoms do not interact directly with one another [20, 27, 28]. Our motivation is to better understand the interplay between interacting Rydberg atoms and a quantized radiation field in view of the physical implementation of quantum state-transfer between atoms and photons for future quantum information processing [34, 36, 37, 29, 30, 31] and photon-photon quantum gates [68, 69, 70, 42]. In Section 2 we formulate the Hamiltonian of the model and discuss its eigenstates and eigenvalues. We then discuss the dynamics of the system for different initial conditions for number (Fock) states of the radiation field (Section 3) and a quantum coherent field (Section 4). In Section 5, we discuss different types of entanglements between various components of the system. In Section 6, we summarize our results and compare the entanglements for a system with Rydberg-interacting atoms with a system of non-interacting atoms corresponding to the Tavis-Cummings model. Detailed derivations of some results discussed in the main text are left to A - D.
@@ 2 Hamiltonian and eigenstates
In this section, we extend the Jaynes-Cummings model for a single atom  [^2] The Jaynes-Cummings model for a single atom is discussed in many textbooks on quantum optics; see, e.g., [71, 72], and the recent review in [13] for a comprehensive list of references. to two Rydberg atoms interacting by the Rydberg blockade. We consider two two-level atoms $a$ and $b$ with ground states |$\ket{g_{a}}$, |$\ket{g_{b}}$ and Rydberg states |$\ket{r_{a}}$, |$\ket{r_{b}}$  [^3] When there is no danger of confusion we omit the labels $a$ and $b$ and denote the ground and Rydberg states for both atoms by |$\ket{g}$ and |$\ket{r}$, respectively., respectively, interacting with a single mode of a quantized radiation field ($f$) with (angular) frequency $\omega_{f}$  [^4] If not stated otherwise, we use the term “frequency” for the angular frequency in units of $\text{rad}\cdot\text{s}^{-1}$.. We assume that the unperturbed transition frequencies $\omega_{0}=\omega_{r}-\omega_{g}$ of the atomic transition |$\ket{g}\leftrightarrow\ket{r}$ of both atoms are equal. The combined system of the two atoms and the quantized field is spanned by product states of the form
@@
We assume that the Rydberg blockade between the atoms is in full effect, so that the doubly excited state, |$\ket{r,r,\phi_{f}}$, is not accessible to the system. Thus, accessible states are spanned by states of the form |$\ket{g,g,n}$, |$\ket{g,r,n}$, and |$\ket{r,g,n}$, with the photon number (Fock) states |$\ket{n}$, $n=0,1,2,\ldots$, of the field. We describe the interaction between the atoms and the quantized field by a Jaynes-Cummings-type Hamiltonian in the rotating wave approximation,  [^5] To simplify notation, in (4) we use the $\otimes$ symbol only to separate operators acting on atoms $a$ and $b$, but not for the field operators $\hat{a}$, $\hat{a}^{{\dagger}}$.
$\hat{H}=\hbar\omega_{f}{\hat{a}}^{{\dagger}}{\hat{a}}-\frac{1}{2}\hbar\omega_{0}\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda_{a}\left({\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\right)\otimes{\hat{P}}_{g}^{\,b}\,+{\hat{P}}_{g}^{\,a}\otimes\hbar\lambda_{b}\left({\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\right)\,.$ (4)
The first term in (4) is the energy of the field with the photon number operator ${\hat{n}}={\hat{a}}^{\dagger}{\hat{a}}$, where ${\hat{a}}^{\dagger}$, $\hat{a}$ are bosonic creation and annihilation operators and $\hbar=h/\left(2\pi\right)$ is the reduced Planck constant. The second term in (4) is the energy of the unperturbed atoms $a$, $b$ (upper labels) expressed in terms of the Pauli operators ${\hat{\sigma}}_{z}=\mathinner{|{g}\rangle\langle{g}|}-\mathinner{|{r}\rangle\langle{r}|}$, where the zero of energy is chosen as the average of the energies of |$\ket{g}$ and |$\ket{r}$. The third and fourth terms in (4) describe processes in which either atom $a$ or atom $b$ is excited from |$\ket{g}$ to |$\ket{r}$ by absorption of one photon, or decays from |$\ket{r}$ to |$\ket{g}$ by emission of one photon, with coupling constants $\lambda_{a}$, $\lambda_{b}$ for atoms $a$, $b$, respectively; ${\hat{\sigma}}_{+}=\mathinner{|{r}\rangle\langle{g}|}$ and ${\hat{\sigma}}_{-}=\mathinner{|{g}\rangle\langle{r}|}$ are raising and lowering Pauli operators for atoms $a$, $b$ (upper labels in (4)). In the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ the Pauli operators are represented by $2\times 2$ matrices  [^6] Note that in the basis {|$\left\{{\ket{g},\ket{r}}\right\}$ (in this order) the raising and lowering Pauli matrices $\sigma_{+}$, $\sigma_{-}$ in (5) are defined the opposite way compared to raising and lowering Pauli matrices for spin-$1/2$ particles.
@@
Since the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system due to the Rydberg blockade, excitation and decay of one atom can only occur when the other atom is in its ground state |$\ket{g}$; this effect is incorporated in the Hamiltonian (4) by the projection operators ${\hat{P}}_{g}^{\,a}=\mathinner{|{g_{a}}\rangle\langle{g_{a}}|}$ and ${\hat{P}}_{g}^{\,b}=\mathinner{|{g_{b}}\rangle\langle{g_{b}}|}$ to the ground states of atoms $a$ and $b$, respectively. The fact that the doubly excited state |$\ket{r,r,\phi_{f}}$ is not accessible to the system creates a direct interaction between the atoms in the Hamiltonian (4), in addition to the interaction of the atoms with the quantized radiation field. This makes our model fundamentally different from previous studies using the Dicke model and Tavis-Cummings model, in which the projection operators ${\hat{P}}_{g}^{\,a}$ and ${\hat{P}}_{g}^{\,b}$ are absent and hence the atoms do not interact directly with one another [20, 27, 28, 29] (see Section 6).
@@
$\hat{H}=\hbar\omega_{f}\hat{N}+\frac{1}{2}\hbar\Delta\left({\hat{\sigma}}_{z}^{a}+{\hat{\sigma}}_{z}^{b}\right)\,+\hbar\lambda\Big{[}\Big{(}{\hat{\sigma}}_{+}^{a}{\hat{a}}+{\hat{\sigma}}_{-}^{a}{\hat{a}}^{{\dagger}}\Big{)}\otimes{\hat{P}}_{g}^{\,b}\,+{\hat{P}}_{g}^{\,a}\otimes\Big{(}{\hat{\sigma}}_{+}^{b}{\hat{a}}+{\hat{\sigma}}_{-}^{b}{\hat{a}}^{{\dagger}}\Big{)}\Big{]}\,,$ (9)
@@
|$\ket{\psi_{1}^{(n)}}:=\ket{r,g,n}\,,\,\,\ket{\psi_{2}^{(n)}}:=\ket{g,r,n}\,,\,\,\ket{\psi_{3}^{(n)}}:=\ket{g,g,n+1}\,.$ (11)
Thus, the total Hilbert space $\mathbb{H}$ of the system decouples into 3-dimensional subspaces $\mathbb{H}_{n}=\text{span}\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$, and $\hat{H}$ in (9) can be expressed as a direct sum ${\hat{H}}=\mathop{\oplus}\limits_{n=0}^{\infty}{\hat{H}}^{(n)}$ where the action of each component ${\hat{H}}^{(n)}$ is confined to the subspace $\mathbb{H}_{n}$  [^8] Without a direct interaction between the atoms as in the Tavis-Cummings model, the doubly excited state |$\ket{r,r,\phi_{f}}$ is accessible to the system. As a result, the system decouples into 4-dimensional subspaces (rather than 3-dimensional) that include the state |$\ket{r,r,n-1}$ (for $n\geq 1$) in addition to the 3 states in (11). See Section 6, D, and the text in the second paragraph below equation (1) in [28].. In the basis {|$\{\ket{\psi_{1}^{(n)}},\ket{\psi_{2}^{(n)}},\ket{\psi_{3}^{(n)}}\}$ the Hamiltonian ${\hat{H}}^{(n)}$ is represented by a $3\times 3$ matrix
@@
|$\ket{\text{asym},n}:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}-\ket{\psi_{2}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}-\ket{g,r,n}\right)\,.$ (17)
@@
is the $\Delta$-dependent Rabi frequency for the two-atom system. The normalization constants in (16) are
$N_{\pm}=\sqrt{\left(\Omega_{n}\pm\Delta/2\right)^{2}+2\lambda^{2}(n+1)}$.
@@
Geometrically, $R$ in (22) describes a rotation of the eigenvectors |±$\ket{\pm,n}$ by the angle $\varphi_{n}(\Delta)$ about the fixed axis given by the $\Delta$-independent eigenvector |$\ket{\text{asym},n}$ (see Figure 3).
@@
|$\ket{\psi_{i}^{(n)}}=R_{\text{asym},\,i}^{T}\ket{\text{asym},n}+R_{+,\,i}^{T}\ket{+,n}+R_{-,\,i}^{T}\ket{-,n}\,,\,\,i=1,2,3\,,$ (26)
@@
|±$\ket{\pm,n}=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{\psi_{3}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{\text{sym},n}\pm\ket{g,g,n+1}\right)\,,$ (34)
@@
|$\ket{\text{sym},n}:=\frac{1}{\sqrt{2}}\left(\ket{\psi_{1}^{(n)}}+\ket{\psi_{2}^{(n)}}\right)=\frac{1}{\sqrt{2}}\left(\ket{r,g,n}+\ket{g,r,n}\right)\,.$ (35)